# A Closer Look at Time Steps is Worthy of Triple Speed-Up for Diffusion Model Training

### v2

@@ Appendix A More Detail of Experiments
A.1 Statement
Experimental design note: Referring to previous works (Peebles & Xie, 2023; Go et al., 2023; Park et al., 2024b; Choi et al., 2022; Hang et al., 2023; Esser et al., 2024; Xu et al., 2024), our experiments train 50K, 100K, and 400K in MetFaces 256 × 256, FFHQ 256 × 256, and ImageNet-1K, respectively. we know that DiT used 7M iterations to train ImageNet in the original work Peebles & Xie (2023), which 1) is not the amount of resource usage that can be achieved in general research. Meanwhile, 2) the focus of this paper is on the acceleration effect of training, and a direct comparison of the ultimate convergence stage is very unnecessary for the research topic of this paper.
A.2 Instructions for code
Code is provided in the link at the very beginning of the main paper.
A.3 Architecture and Training Recipe.
@@ A.4 Datasets
A.5 Detail of MDT + SpeeD Experiment
@@
Following the official implementation of MDT, We utilize DiT-S/2 and MDT-S/2 as our base architecture, whose total block number both is 12 and the number of decoder layers in MDT is $2$. We employ the AdamW (Loshchilov & Hutter, 2017) optimizer with constant learning rate 1e-4 using $256$ batch size without weight decay on class-conditional ImageNet with an image resolution of $256^{2}$. We perform training on the class-conditional ImageNet dataset with images of resolution 256x256. The diffusion models are trained for a total of 1000K iterations, utilizing a mask ratio of 0.3.
A.6 Detail of FDM + SpeeD Experiment
FDM add the momentum to the forward diffusion process with a scale that control the weight of momentum for faster convergence to the target distribution. Following official implementation, we train diffusion models of EDM and FDM. We retrain these official network architecture which is U-Net with positional time embedding with dropout rate 0.13 in training. We adopt Adam optimizer with learning rate 1e-3 and batch size 512 to train each model by a total of 200 million images of $32^{2}$ CIFAR-10 dataset. During training, we adopt a learning rate ramp-up duration of 10 Mimgs and set the EMA half-life as 0.5 Mimgs. For evaluation, EMA models generate 50K images using EDM sampler based on Heun’s $2^{nd}$ order method (Süli & Mayers, 2003).
A.7 Text-to-Image Experiment Detail
@@
A.8 Theoretical Analysis
A.8.1 Notations
@@
A.8.2 Auxiliary Lemma and Core Theorem
@@
In DDPM (Ho et al., 2020), using a simple equivariant series $\{\beta_{t}\}_{t\in[T]}$ to simplify the complex cumulative products $\{\bar{\alpha}_{t}\}_{t\in[T]}$, we obtain the following auxiliary upper bound of $\bar{alpha}_{t}$.
@@ B.1 Additional Experiments
B.1.1 100K iterations cut on FFHQ
A cut of training process comparison on FFHQ dataset are shown in Tab. 10.
Schedules linear quadratic cosine
DiT-XL/2 (Peebles & Xie, 2023) 7.8 7.6 8.0
ANT-UW (Go et al., 2023) 6.3 7.1 8.3
TskDif-Cur (Park et al., 2024b) 6.7 6.7 7.2
P2 (Choi et al., 2022) 10.1 10.2 12.3
MinSNR (Hang et al., 2023) 7.6 7.6 11.2
Log-Normal (Esser et al., 2024) 15.6 14.9 19.2
CLTS (Xu et al., 2024) 7.9 7.7 9.1
SpeeD(ours) 5.8 5.8 6.4
Table 10: 100K iterations cut on FFHQ.
B.1.2 Guidance Scale
Ablation experiments about guidance scale in related conditional generation task are shown in Tab. 11.
Guidance 0.75 1 1.5 2 3
t2i DiT-XL/2 20.1 22.5 27.4 31.0 36.2
t2i SpeeD 19.4 20.8 25.3 28.8 34.8
ccg EDM 1.8 1.8 1.9 1.9 2.5
ccg SpeeD on EDM 1.7 1.7 1.8 1.8 2.3
Table 11: Text-to-image (t2i) task on MS-COCO at 400K iteration and EDM on CIFAR-10 with 200M training images for class-conditional generation (ccg).
B.1.3 DiT Architecture
Ablation experiments about DiTs of different architectures are shown in Tab. 12.
S/2 B/2 XL/2
DiTs 18.1 12.9 7.8
SpeeD 15.3 10.8 5.8
Table 12: Different size of DiT with 100K iterations on FFHQ.
B.1.4 Detailed Training Process
@@ Appendix C More Related Works
We discuss other works related to SpeeD, including Text to Image and Video generation. Another point to mention is that we learn from InfoBatch (Qin et al., 2023) in writing.
Text to image generation with diffusion models Text-to-image generation has emerged as a hotly contested and rapidly evolving field in recent years, with an explosion of related industrial products springing up (Saharia et al., 2022; Rombach et al., 2022; Betker et al., 2023; Chen et al., 2023a; Esser et al., 2024). Convert textual descriptions into corresponding visual content, models not only learn to synthesize image content but also ensuring alignment with the accompanying textual descriptions. To better align images with textual prompt guidance, previous work has primarily focused on enhancements in several schemes including strengthening the capacity of text encoder (Raffel et al., 2020; Radford et al., 2021) improving the condition plugin module in diffusion model (Zhang et al., 2023b), improving data quality (Betker et al., 2023).
Video generation with diffusion models. As diffusion models achieve tremendous success in image generation, video generation has also experienced significant breakthroughs, marking the field’s evolution and growth. Inspired by image diffusion, pioneering works such as RVD Yang et al. (2022) and VDM Ho et al. (2022b) explore video generation using diffusion methods. Utilizing temporal attention and latent modeling mechanisms, video diffusion has advanced in terms of generation quality, controllability, and efficiency Ho et al. (2022a); Singer et al. (2022); Zhou et al. (2022); He et al. (2022); Zhang et al. (2023a); Guo et al. (2023); Wu et al. (2022); Wang et al. (2023). Notably, Stable Video Diffusion Blattmann et al. (2023) and Sora Brooks et al. (2024) achieve some of the most appealing results in the field.
Other diffusion acceleration works To achieve better results with fewer NFE steps, Consistency Models Song et al. (2023) and Consistency Trajectory Models Kim et al. (2023) employ consistency loss and novel training methods. Rectified Flow Liu et al. (2022), followed by Instaflow Liu et al. (2023), introduces a new perspective to obtain straight ODE paths with enhanced noise schedule and improved prediction targets, together with the reflow operation.
@@ Appendix D Visualization
Figure 8: Comaprisons to the baseline and other acceleration methods.
Comparisons to the baseline and other methods.
The Fig. 8 compares our method (Ours) with the baseline and other acceleration methods (P2 and Min-SNR) in terms of FID scores at various training iterations (K). The baseline method starts with a high FID score of 335.2 and gradually decreases to 12.8, showing slow convergence and less sharp final images. P2 begins with a slightly higher FID score of 357.9 and reduces to 15.0, improving faster than the baseline but still exhibiting slower convergence compared to Min-SNR and our method. Min-SNR starts with 334.1 and achieves a final score of 12.2, producing clearer and higher quality images consistently compared to the baseline and P2. Our method stands out, starting with a significantly lower FID score of 289.3 and rapidly dropping to 9.9, demonstrating the fastest convergence and the best overall image quality. This rapid decrease highlights the efficiency of our approach, requiring fewer iterations to produce high-quality results. Overall, our method outperforms the baseline, P2, and Min-SNR in both speed and quality of image generation, proving its effectiveness in accelerating the training of diffusion models.

### v3

@@ Appendix A More Detail of Experiments
A.1 Architecture and Training Recipe.
@@ A.2 Datasets
A.3 Detail of MDT + SpeeD Experiment
@@
Following the official implementation of MDT, We utilize DiT-S/2 and MDT-S/2 as our base architecture, whose total block number both is 12 and the number of decoder layers in MDT is $2$. We employ the AdamW [39] optimizer with constant learning rate 1e-4 using $256$ batch size without weight decay on class-conditional ImageNet with an image resolution of $256^{2}$. We perform training on the class-conditional ImageNet dataset with images of resolution 256x256. The diffusion models are trained for a total of 1000K iterations, utilizing a mask ratio of 0.3.
Table 8: The ingredients of generalized curves $\dot{\Delta}$ and $\dot{\Sigma}$ schedules about mainstream SDE designs, including VP, VE [59], EDM [28].
      Schedules       $s$       $\sigma^{2}$       $\dot{s}$       $\dot{\sigma}$
      VE       $1$       $t$       00       $1$
      EDM       $1$       $t^{2}$       00       $2t$
A.4 Detail of FDM + SpeeD Experiment
FDM add the momentum to the forward diffusion process with a scale that control the weight of momentum for faster convergence to the target distribution. Following official implementation, we train diffusion models of EDM and FDM. We retrain these official network architecture which is U-Net with positional time embedding with dropout rate 0.13 in training. We adopt Adam optimizer with learning rate 1e-3 and batch size 512 to train each model by a total of 200 million images of $32^{2}$ CIFAR-10 dataset. During training, we adopt a learning rate ramp-up duration of 10 Mimgs and set the EMA half-life as 0.5 Mimgs. For evaluation, EMA models generate 50K images using EDM sampler based on Heun’s $2^{nd}$ order method [61].
A.5 Text-to-Image Experiment Detail
@@
A.6 Theoretical Analysis
A.6.1 Notations
@@
A.6.2 Auxiliary Lemma and Core Theorem
@@
In DDPM [20], using a simple equivariant series $\{\beta_{t}\}_{t\in[T]}$ to simplify the complex cumulative products $\{\bar{\alpha}_{t}\}_{t\in[T]}$, we obtain the following auxiliary upper bound of $\bar{alpha}_{t}$.
@@ B.1 Additional Experiments
Experimetns on DiT-S/8.
Further comparison on DiT of smaller scales are reported in Tab. 10. The datasets include ImageNet-1K and Celeb-A.
ImageNet-1K (FID $\downarrow$)
#steps DiT-S/8 P2 Min-SNR Log-normal CLTS ours
10K 399.9 398.6 398.4 399.7 399.6 400.8
20K 380.0 365.0 368.0 387.5 381.9 379.2
40K 200.0 207.6 208.6 365.9 231.6 191.5
Celeb-A (FID $\downarrow$)
10K 408.7 412.5 408.7 410.4 408.0 407.8
20K 386.7 366.8 386.7 394.0 386.4 377.9
40K 271.6 271.1 271.6 293.0 258.8 254.9
Celeb-A (IS $\uparrow$)
10K 1.50 1.49 1.50 1.50 1.50 1.50
20K 1.49 1.48 1.49 1.50 1.49 1.63
40K 3.29 3.70 3.29 2.55 3.46 3.87
Table 10: Comparison on ImageNet-1K and Celeb-A. FID and IS are reported. The baseline is measured on DiT-S/8, with global batchsize of 16. Other settings are default.
More baslines.
Comparison between SpeeD and BS [78] and B-TTDM [79] are shown in Tab. 11.
Detailed ablation study.
FID-10K on 10K/20K/40K/50K iterations are provided in Tab. 12 the model is DiT-S/8 with batchsize of 16.
#steps BS [78] B-TTDM [79] ours
100K 155.9 157.4 155.2
200K 152.2 150.9 151.7
400K 141.8 140.5 139.2
Table 11: New baselines to be added. The settings follow BS. (FID)
#steps 10K 20K 40K 50K
DiT-S/8 399.9 380.0 200.0
$\lambda=0.5$ 400.7 376.7 207.4 202.1
$\lambda=0.6$ 400.8 379.2 191.5 191.2
$\lambda=0.8$ 400.3 379.2 203.2 200.5
$\tau=600$ 401.0 382.6 210.1 198.5
$\tau=700$ 400.8 379.2 191.5 191.2
$\tau=800$ 399.5 380.9 200.1 200.0
$k=1$ 400.4 388.0 214.5 202.4
$k=2$ 400.8 379.2 191.5 191.2
$k=10$ 400.4 380.5 231.7 206.7
Table 12: Detailed ablation across training steps. FID-10K is reported. DiT-S/8 serves as the baseline. Global batchsize is 16.
B.1.1 Detailed Training Process
@@ Appendix C More Related Works
We discuss other works related to SpeeD, including Text to Image and Video generation. Another point to mention is that we learn from InfoBatch [48] in writing.
Text to image generation with diffusion models Text-to-image generation has emerged as a hotly contested and rapidly evolving field in recent years, with an explosion of related industrial products springing up [54, 51, 2, 6, 11]. Convert textual descriptions into corresponding visual content, models not only learn to synthesize image content but also ensuring alignment with the accompanying textual descriptions. To better align images with textual prompt guidance, previous work has primarily focused on enhancements in several schemes including strengthening the capacity of text encoder [50, 49] improving the condition plugin module in diffusion model [71], improving data quality [2].
Video generation with diffusion models. As diffusion models achieve tremendous success in image generation, video generation has also experienced significant breakthroughs, marking the field’s evolution and growth. Inspired by image diffusion, pioneering works such as RVD [68] and VDM [22] explore video generation using diffusion methods. Utilizing temporal attention and latent modeling mechanisms, video diffusion has advanced in terms of generation quality, controllability, and efficiency [21, 56, 81, 18, 70, 15, 65, 63]. Notably, Stable Video Diffusion [3] and Sora [5] achieve some of the most appealing results in the field.
Other diffusion acceleration works To achieve better results with fewer NFE steps, Consistency Models [60] and Consistency Trajectory Models [29] employ consistency loss and novel training methods. Rectified Flow [36], followed by Instaflow [37], introduces a new perspective to obtain straight ODE paths with enhanced noise schedule and improved prediction targets, together with the reflow operation. DyDiT [73] incorporates dynamic neural networks [16, 73, 72] into diffusion models, achieving significant acceleration.